OpenStudy (anonymous):

34. Find the 6th term of a geometric sequence with t1=5 and r =-1/2

ganeshie8 (ganeshie8):

\(t_n = ar^{n-1} \) a = first term

ganeshie8 (ganeshie8):

\(t_6 = 5(\frac{-1}{2})^{6-1}\) =\(5(\frac{-1}{2})^{5}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

-5/32

ganeshie8 (ganeshie8):

Yes !